OpenStudy (anonymous):

The cash price of a washing machine is £470. Mrs Dan buys the washing machine. She pays a deposit of 30% of the cash price and £30.25 each month for 12month. Mrs dan pays, in total, more than £470. How much more?

OpenStudy (anonymous):

Total payment = deposit + monthly payment t = 30% * 470 + 12 * 30.25 = 81 + 363 = £444 Left over: 470 - 444 = £26